# When Measurement Mediates the Effect of Interest

### v2

@@ 4.2 Estimators Compared
We compared the finite sample performance of different approaches within the Two-Stage framework. For identification and estimation of the cluster-level outcome $Y^{c*}=\mathbb{P}(Y2=1\mid Y1^{*}=1)$ in Stage 1, we considered two approaches aiming to directly estimate this conditional probability in each cluster. First, we considered the empirical mean outcome among those measured: ^$\hat{\mathbb{P}}(Y2=1\mid\Delta=1)$. We refer to this approach as “Screened”. Second, we considered the empirical mean outcome among persons known to be in the target population and, thus, eligible for the outcome: ^$\hat{\mathbb{P}}(Y2=1\mid Y1=1)$. We refer to this approach as “Eligible”. While these estimators are intuitive, they do not estimate the cluster-level endpoint $Y^{c*}$ and are expected to be biased for the total effect, even under MCAR (Appendix A).
For Stage 1, we also considered two approaches based on re-expressing the cluster-level outcome $Y^{c*}$ as the ratio: $\mathbb{P}(Y2=1,Y1^{*}=1)\div\mathbb{P}(Y1^{*}=1)$. In each, the numerator was estimated by taking the empirical proportion with the outcome: ^$\hat{\mathbb{P}}(Y2=1)$. For the denominator, we considered the MCAR assumption and the unadjusted estimator, calculated as the empirical proportion who are known to be in the target population: ^$\hat{\mathbb{P}}(Y1=1\mid\Delta=1)$. For the denominator, we also considered the MAR assumption for identification and used TMLE to estimate $\mathbb{E}[\mathbb{E}(Y1\mid\Delta=1,W)]$, adjusting for differences in covariate values $W$ between persons screened versus not screened. Within TMLE, we employed Super Learner using 10-fold cross-validation to combine estimates from GLMs, multivariate adaptive regression splines, and the mean. For these two ratio-based approaches, we divided the estimated numerator by the estimated denominator to obtain the cluster-level endpoint estimate $\hat{Y}^{c}$.
When evaluating the total effect in Stage 2, we implemented the unadjusted effect estimator for all approaches. Additionally for TMLE, we applied Adaptive Pre-specification using 5-fold cross validation to select among working GLMs adjusting for at most 1 cluster-level covariate. For all estimators inference was based on the estimated influence curve. With 1000 iterations, we compared performance with the metrics of bias, average standard error estimate, Monte Carlo standard deviation, 95% confidence interval coverage, and power attained.
@@ 4.3 Results
Table 1 summarizes the estimator performance when there was total effect of 2.24%. As expected, the two direct approaches to Stage 1 were the most biased and had the lowest 95% confidence interval coverage. Specifically, the Screened approach estimating $Y^{c}$ with ^$\hat{\mathbb{P}}(Y2=1\mid\Delta=1)$ targeted the wrong parameter, underestimated the total effect by 2.07% on average, and only achieved 47.4% confidence interval coverage. Likewise, the Eligible approach estimating $Y^{c}$ with ^$\hat{\mathbb{P}}(Y2=1\mid Y1=1)$ also targeted the wrong parameter, also underestimated the total effect (bias=-1.72%), but achieved a higher coverage (88.7%) due to being a more variable estimator.
Table 1: Results from 1000 simulated trials with 150 clusters and a true effect of $2.24\%$
Method Pt ($95\%$ CI) Bias $\hat{\sigma}$ $\sigma$ Cover. Power
Stage 1 Stage 2
Screened Unadjusted 0.17(-1.82 , 2.17) -2.07 0.0102 0.0104 47.4 6.0
Eligible Unadjusted 0.52 (-4.53 , 5.60) -1.72 0.0258 0.0265 88.7 6.7
Unadjusted Unadjusted 3.11 (-0.14 , 6.38) 0.87 0.0166 0.0170 90.8 45.2
TMLE Unadjusted 2.17 (-0.63 , 4.97) 0.07 0.0143 0.0144 94.7 33.5
TMLE TMLE 2.16 (0.42 , 3.89) -0.08 0.0088 0.0086 95.2 68.4
@@
  • Cover.: Proportion of 95% confidence interval containing the true effect (in %)
@@
Performance improved with the ratio-based approaches to Stage 1. However, due to relying on the MCAR assumption, the approach using the unadjusted estimator in Stage 1 was still biased. It over-estimated the total effect by 0.87% on average and resulted in poor confidence interval coverage of 90.8%. In contrast, relaxing the missing data assumption and using TMLE for flexible estimation in Stage 1 resulted in minimal bias and nominal 95%confidence interval coverage. Crucially, using TMLE in Stage 2 was much more efficient, doubling the achieved power (68.4%) as compared using the unadjusted estimator in Stage 2 (33.5%).
@@ 5 Extensions
We now present an extension of the above methods to (1) accommodate time-varying covariates and (2) allow the exposure to impact the population of interest. Figure 3 provides an updated causal graph, where $\bm{L}$ denotes the matrix of post-baseline, individual-level covariates. In OPAL, these covariates could include measures of mobility and alcohol use. As shown in Figure 3, there are now several pathways of the intervention effect. As before, the intervention $A^{c}$ directly impacts outcomes $\bm{Y2}$ and indirectly through its effect on health screening $\bm{\Delta}$. Additionally, the intervention now indirectly impacts outcomes through its effect on the time-varying covariates $\bm{L}$ and on underlying status defining who is in the target population $\bm{Y1}^{*}$. Importantly, the time-varying covariates $\bm{L}$ effectively confound the relationship between health screening $\bm{\Delta}$ and health outcomes $\bm{Y2}$. The corresponding non-parametric structural equations are provided in Appendix C.
@@
Our interest remains in the total effect $\Psi^{*}=\mathbb{E}[Y^{c*}(1)-Y^{c*}(0)]$. However, the target population is now a function of the intervention: $Y^{c*}(a^{c})=\mathbb{P}[Y2(a^{c})=1\mid Y1^{*}(a^{c})=1]$. Despite a more complicated causal model and causal estimand, the Two-Stage approach is still applicable. Stratification on each cluster in Stage 1 obviates the multilevel-mediation-missing data problem as well as the missing-data equivalent to time-dependent confounding [27].
Our Stage 1 causal estimand remains the same and is the counterfactual probability of the outcome for persons in the underlying target population:
@@
As before, the numerator is trivially identified as the proportion with the outcome: $\mathbb{P}(Y2=1)$. Now, however, we can account for baseline and time-varying covariates when identifying and estimating the denominator.
Specifically, we assume that within values of baseline and post-baseline covariates, health status among those screened is representative among health status among those missed: $Y1^{*}\perp\Delta\mid W,L$. Additionally, we require that the probability of being measured within covariate strata is bounded away from 0: $P(\Delta=1\mid W=w,L=l)>0\ \text{ for all possible }w,l$. Under these assumptions, we can write the statistical estimand for the denominator as $\mathbb{E}[\mathbb{E}(Y1\mid\Delta=1,W,L)]$. Estimation proceeds as before using TMLE. Then we combine the numerator and denominator estimates to generate $\hat{Y}^{c}$, repeat for all $J$ clusters, and conduct Stage 2 estimation and inference using TMLE with Adaptive Pre-specification.
@@ 6 Discussion
Novel challenges and opportunities arise when health promotion strategies aim to expand reach into the target population and outcomes among those reached. Specifically, the strategy improves health outcomes directly and indirectly through expanded reach, while outcomes are only measured among those reached. In this common scenario, existing methods to address differential outcome measurement would effectively block the indirect effect and bias the total effect towards the null. Using causal models, we specified the data generating process and defined the causal estimand with Counterfactual Strata Effects to capture the intervention effect in the underlying target population, without enforcing 100% compliance on measurement. Then we extended the Two-Stage TMLE framework for estimation and inference. Our simulations demonstrated that our proposed approach obtained nominal confidence interval coverage and more precision than more traditional approaches. We plan to use our novel approach in the primary, pre-specified analysis of the OPAL trial, which is evaluating the impact of community-based strategies use of biomedical HIV prevention among adults at drinking venues in Kenya and Uganda.
@@
This work has limitations, and future work is needed. First, the Two-Stage TMLE framework obviates the multilevel-mediation-missing data problem by fully stratifying on each cluster in Stage 1. This can lead to poor data support when there are small clusters and/or a large adjustment set. We are exploring strategies to adaptively pool across clusters within arm in Stage 1. Second, to match the motivating example, we focused on a binary outcome. However, our methods are equally applicable to other types of outcomes. Also to match the motivating example, we studied cluster randomized trials, which randomize groups to the intervention and comparator strategy. We plan to extend this work to observational studies where the group-level strategy is delivered in a non-randomized strategy. In such a setting, we would need to adjust for cluster-level confounding in Stage 2 of Two-Stage TMLE. We also plan to extend this work to individually randomized trials where the health promotion strategy is delivered at the individual-level. Two-Stage TMLE, designed for hierarchical data, would not be immediately applicable in such a setting; however, we hypothesize that creating artificial clusters within arm and implementing the Two-Stage procedure (with appropriate weights) should facilitate estimation of the total effect, despite the mediation-missing data problem. Future work is warranted.

### v3

@@ 4.2 Estimators Compared
For Single-Stage estimators, we implemented GEE among those who screened ($\Delta=1$), adjusting for individual-level and cluster-level covariates ($W1,W2,W3,E1^{c},E2^{c}$), and accounting for clustering with an independent working correlation matrix. To do so, we used the MRStdCRT R package to target the marginal effect on the absolute scale [29].
We also implemented Single-Stage TMLE among participants known to be in the focus population ($Y1=1$), adjusting for the same covariates, and accounting for clustering by generating a cluster-level influence curve [53]. To do so, we used the ltmle R package with id to indicate cluster membership [54]. While these Single-Stage approaches are more traditional, they condition on the mediator $(\Delta=1)$ or a descendant of the mediator ($Y1=1$) and expected to be biased for the total effect in the underlying focus population.
Within the Two-Stage framework, we evaluated various approaches for identification and estimation of the cluster-level outcome $Y^{c*}=\mathbb{P}(Y2=1\mid Y1^{*}=1)$ in Stage 1. First, we considered the empirical mean outcome among those measured: $\hat{\mathbb{P}}(Y2=1\mid\Delta=1)$. We refer to this approach as “Screened”. Second, we considered the empirical mean outcome among persons known to be in the focus population and, thus, eligible for the outcome: $\hat{\mathbb{P}}(Y2=1\mid Y1=1)$. We refer to this approach as “Eligible”. While these estimators are intuitive, they do not estimate the cluster-level endpoint $Y^{c*}$ and are expected to be biased for the total effect, even under MCAR (Appendix A of the Supporting Information). For Stage 1, we also considered two approaches based on re-expressing the cluster-level outcome $Y^{c*}$ as the ratio: $\mathbb{P}(Y2=1,Y1^{*}=1)\div\mathbb{P}(Y1^{*}=1)$. In each, the numerator was estimated by taking the empirical proportion with the outcome: $\hat{\mathbb{P}}(Y2=1)$. For the denominator, we considered the MCAR assumption and the unadjusted estimator, calculated as the empirical proportion who are known to be in the focus population: $\hat{\mathbb{P}}(Y1=1\mid\Delta=1)$. For the denominator, we also considered the MAR assumption for identification and used TMLE to estimate $\mathbb{E}[\mathbb{E}(Y1\mid\Delta=1,W)]$, adjusting for differences in covariate values $W$ between persons screened versus not screened. For these two ratio-based approaches, we divided the estimated numerator by the estimated denominator to obtain the cluster-level endpoint estimate $\hat{Y}^{c}$. When evaluating the total effect in Stage 2, we implemented the unadjusted effect estimator for all approaches. Specifically, we took the difference in the empirical mean of the cluster-level endpoints $\hat{Y}^{c}$ by arm. For the approach using TMLE in Stage 1, we implemented TMLE with Adaptive Pre-specification in Stage 2. Specifically, we used cross-validation to select among working GLMs adjusting for at most 1 cluster-level covariate.
@@ 4.3 Results
In these simulations, the true value of the total effect was 4.28%. For $J=\{20,30,50,70\}$ clusters with an average size of 200 persons, Table 1 provides the following metrics across 1000 iterations: the average of the point estimate and 95% confidence intervals, bias (average deviation of point estimates from the true effect), the Monte Carlo standard deviation, the average standard error estimate, coverage (proportion of 95% confidence intervals containing the true effect), and attained power (proportion of tests rejecting the null hypothesis).
As expected, both Single-Stage approaches were meaningfully biased towards the null. Specifically, both GEE and the Single-Stage TMLE underestimated the total effect by conditioning on a mediator (screening) or a descendant of the mediator (eligibility). For both estimators, the resulting confidence interval coverage was much less than the nominal rate of 95% and decreased with increasing sample size. Indeed, at a sample size of $J=70$, coverage was 52.5% for GEE and was $53.5\%$ for the Single-Stage TMLE.
Within the Two-Stage framework, the two approaches estimating the cluster-level endpoint $\hat{Y}^{c}$ among those screened $\hat{\mathbb{P}}(Y2=1\mid\Delta=1)$ and among those eligible $\hat{\mathbb{P}}(Y2=1\mid Y1=1)$ were also meaningfully biased. This was expected, because both did not target the correct estimand in Stage 1 (Appendix A of the Supporting Information). The “Screened” approach was biased toward the null. Its 95% confidence interval coverage ranged from 81.3% for the smallest sample size to 52.4% for the largest sample size. On average, the “Eligible” approach generated a negative effect estimate — suggesting the positive intervention was actually harmful. Its 95% confidence interval coverage ranged from 62.5% for the smallest sample size to 13.5% for the largest sample size.
Table 1: Performance of Single-Stage and Two-Stage approaches across 1000 simulated trials with $J=\{20,30,50,70\}$ clusters and a true effect of 4.28%.
Method J Pt (95% CI) Bias $\hat{\sigma}$ $\sigma$ Coverage Power
Single-Stage Methods
GEE among screened ($\Delta=1$) adjusting for ($W,E^{c}$) 20 0.89(9.61, 11.38)0.89\;(-9.61,\;11.38) -3.40 0.0411 0.0502 91.5 3.3
30 0.74(6.10, 7.59)0.74\;(-6.10,\;7.59) -3.54 0.0300 0.0335 83.6 6.2
50 0.89(3.73, 5.50)0.89\;(-3.73,\;5.50) -3.40 0.0212 0.0230 69.7 5.5
70 0.73(2.92, 4.37)0.73\;(-2.92,\;4.37) -3.56 0.0174 0.0183 52.2 5.9
TMLE among eligible ($Y1=1$) adjusting for ($W,E^{c}$) 20 1.58(3.51, 6.67)1.58\;(-3.51,\;6.67) -2.71 0.0285 0.0243 78.9 14.1
30 1.56(2.74, 5.85)1.56\;(-2.74,\;5.85) -2.73 0.0234 0.0210 72.2 12.8
50 1.67(1.63, 4.98)1.67\;(-1.63,\;4.98) -2.61 0.0174 0.0164 66.6 16.9
70 1.57(1.21, 4.35)1.57\;(-1.21,\;4.35) -2.71 0.0141 0.0139 53.5 19.6
Two-Stage Methods
“Screened” in Stage1 & Unadjusted for Stage2 20 0.48(6.82, 7.77)0.48\;(-6.82,\;7.77) -3.81 0.0356 0.0347 81.3 5.6
30 0.69(5.22, 6.59)0.69\;(-5.22,\;6.59) -3.60 0.0303 0.0288 76.6 6.4
50 0.78(3.62, 5.17)0.78\;(-3.62,\;5.17) -3.50 0.0224 0.0224 64.0 6.5
70 0.70(3.03, 4.44)0.70\;(-3.03,\;4.44) -3.58 0.0194 0.0191 52.4 7.9
“Eligible” in Stage1 & Unadjusted for Stage2 20 1.86(9.30, 5.58)-1.86\;(-9.30,\;5.58) -6.14 0.0363 0.0354 62.5 8.5
30 1.77(7.78, 4.25)-1.77\;(-7.78,\;4.25) -6.05 0.0309 0.0294 48.8 9.6
50 1.65(6.17, 2.88)-1.65\;(-6.17,\;2.88) -5.93 0.0239 0.0231 28.7 12.7
70 1.73(5.55, 2.10)-1.73\;(-5.55,\;2.10) -6.01 0.0192 0.0195 13.5 14.0
Unadjusted ratio in Stage1 & Unadjusted for Stage2 20 2.59(2.00, 7.17)2.59\;(-2.00,\;7.17) -1.70 0.0222 0.0218 89.3 20.4
30 2.68(1.00, 6.36)2.68\;(-1.00,\;6.36) -1.60 0.0195 0.0180 83.7 30.9
50 2.81(0.03, 5.59)2.81\;(0.03,\;5.59) -1.47 0.0144 0.0142 81.3 50.0
70 2.73(0.38, 5.08)2.73\;(0.38,\;5.08) -1.56 0.0120 0.0120 74.8 62.2
TMLE ratio in Stage1 & Unadjusted for Stage2 20 4.10(0.74, 8.93)4.10\;(-0.74,\;8.93) -0.19 0.0236 0.0230 94.5 37.7
30 4.21(0.32, 8.09)4.21\;(0.32,\;8.09) -0.08 0.0205 0.0190 93.6 55.4
50 4.32(1.39, 7.26)4.32\;(1.39,\;7.26) 0.04 0.0151 0.0150 94.7 81.8
70 4.26(1.78, 6.73)4.26\;(1.78,\;6.73) -0.03 0.0126 0.0126 94.0 92.3
TMLE ratio in Stage1 & TMLE for Stage2 20 4.16(0.63, 8.94)4.16\;(-0.63,\;8.94) -0.13 0.0216 0.0228 96.2 40.5
30 4.17(0.54, 7.81)4.17\;(0.54,\;7.81) -0.11 0.0171 0.0177 96.0 60.8
50 4.30(1.65, 6.95)4.30\;(1.65,\;6.95) 0.01 0.0126 0.0135 95.6 90.1
70 4.23(2.05, 6.41)4.23\;(2.05,\;6.41) -0.05 0.0108 0.0111 95.4 96.7
@@
  • Coverage: Proportion of 95% confidence interval containing the true effect (in %)
@@
  • “Screened”: $\hat{Y}^{c}=\hat{\mathbb{P}}(Y2=1\mid\Delta=1)$
  • “Eligible”: $\hat{Y}^{c}=\hat{\mathbb{P}}(Y2=1\mid Y1=1)$
  • Unadjusted ratio: $\hat{Y}^{c}=\hat{\mathbb{P}}(Y2=1)$ $\div$ $\hat{\mathbb{P}}(Y1=1\mid\Delta=1)$
  • TMLE ratio: $\hat{Y}^{c}=\hat{\mathbb{P}}(Y2=1)$ $\div$ $\hat{\mathbb{E}}[\hat{\mathbb{E}}(Y1\mid\Delta=1,W)]$
As expected, the performance of the Two-Stage approaches was improved when taking the ratio-based approach to Stage 1. Also as expected, the unadjusted approach to estimating the cluster-level endpoint $\hat{Y}^{c}$ as $\hat{\mathbb{P}}(Y2=1)\div\hat{\mathbb{P}}(Y1=1\mid\Delta=1)$ was still biased, because it failed to account for differences between those screened versus not screened. Its confidence interval coverage remained well-below the target and was in the range of 74.8% to 89.3%. In contrast, relaxing the missing data assumption and using TMLE in Stage 1 resulted in negligible bias and nominal confidence interval coverage across all sample sizes. Additionally, using TMLE in Stage 2 was more efficient, as expected. Indeed, adaptive selection of the cluster-level adjustment covariates resulted in 2.8% to 8.3% higher power, while maintaining 95% confidence interval coverage. Furthermore, as shown in Appendix B of the Supporting Information, the approach maintained strict Type-I error control under the null.
@@ 5 Extensions
We now present an extension of the above methods to (1) accommodate time-varying covariates and (2) allow the exposure to impact the focus population. Figure 3 provides an updated causal graph, where $\bm{L}$ denotes the matrix of post-baseline, individual-level covariates. In OPAL, these covariates could include measures of mobility and alcohol use. As shown in Figure 3, there are now several pathways of the intervention effect. As before, the intervention $A^{c}$ directly impacts outcomes $\bm{Y2}$ and indirectly through its effect on health screening $\bm{\Delta}$. Additionally, the intervention now indirectly impacts outcomes through its effect on the time-varying covariates $\bm{L}$ and on underlying status defining who is in the focus population $\bm{Y1}^{*}$. Importantly, the time-varying covariates $\bm{L}$ effectively confound the relationship between health screening $\bm{\Delta}$ and health outcomes $\bm{Y2}$. The corresponding non-parametric structural equations are provided in Appendix C in the Supporting Information.
@@
Our interest remains in the total effect $\Psi^{*}=\mathbb{E}[Y^{c*}(1)-Y^{c*}(0)]$. However, the focus population is now a function of the intervention strategy: $Y^{c*}(a^{c})=\mathbb{P}[Y2(a^{c})=1\mid Y1^{*}(a^{c})=1]$. Despite a more complicated causal model and causal estimand, the Two-Stage approach is still applicable. In Stage 1, stratification on each cluster obviates the multilevel-mediation-missing data problem as well as the missing-data equivalent to time-dependent confounding [4]. Within each cluster the cluster-level variables $(E^{c},A^{c})$ are fixed. Therefore, our Stage 1 causal estimand remains the same and is the counterfactual probability of the outcome for persons in the underlying focus population:
@@
As before, the numerator is trivially identified as the proportion with the outcome: $\mathbb{P}(Y2=1)$.
Now, however, we can account for baseline and time-varying covariates when identifying and estimating the denominator. Specifically, we assume that within values of baseline and post-baseline covariates, health status among those screened is representative among health status among those missed: $Y1^{*}\perp\Delta\mid W,L$. Additionally, we require that the probability of being measured within covariate values is bounded away from 0: $P(\Delta=1\mid W=w,L=l)>0\ \text{ for all possible }w,l$. Under these assumptions, we can write the statistical estimand for the denominator as $\mathbb{E}[\mathbb{E}(Y1\mid\Delta=1,W,L)]$. Estimation for the denominator proceeds as before using TMLE with a larger adjustment set. Then we combine the numerator and denominator estimates to generate $\hat{Y}^{c}$, repeat for all $J$ clusters, and conduct Stage 2 estimation and inference using TMLE with Adaptive Pre-specification. We emphasize that Stage 1’s identification and estimation of the cluster-level endpoint $Y^{c}=\mathbb{P}(Y2=1)\div\mathbb{E}[\mathbb{E}(Y1\mid\Delta=1,W,L)]$ is done in each cluster separately. This allows the missingness mechanism to vary by cluster and facilitates our adjustment for time-varying covariates impacted by the trial arm.
@@ 6 Discussion
Novel challenges and opportunities arise when health strategies aim to expand reach into the focus population and improve outcomes. Specifically, the strategy improves health outcomes directly and indirectly through expanded reach, while outcomes are only measured among those reached. In this common scenario, existing methods to address differential outcome measurement would effectively block the indirect effect and bias the total effect towards the null. Building on prior work [4], we used causal models to specify the data generating process and define the causal estimand with Counterfactual Strata Effects capturing the intervention effect in the underlying target population, without enforcing 100% compliance on screening. Then we extended the Two-Stage TMLE framework for estimation and inference. Our simulations demonstrated that our proposed approach obtained nominal confidence interval coverage, while more traditional approaches failed. We plan to use our novel approach in the primary, pre-specified analysis of the OPAL trial, which is evaluating the effectiveness of community-based, recruitment strategies on the use of biomedical HIV prevention among adults at alcohol-serving venues in Kenya and Uganda.
@@
This work has limitations, and future work is needed. First, the Two-Stage TMLE framework obviates the multilevel-mediation-missing data problem by fully stratifying on each cluster in Stage 1. This can lead to poor data support when there are small clusters and/or a large adjustment set. We are exploring strategies to adaptively pool across clusters within arm in Stage 1 [37]. We note, however, that pooling approaches would require homogeneity assumptions that are not needed in our Two-Stage framework. Second, to match the motivating example, we focused on a binary outcome. However, our methods are equally applicable to other types of outcomes. Also to match the motivating example, we studied cluster randomized trials, which randomize groups to the intervention and comparator strategy. We plan to extend this work to observational studies where the group-level strategy is delivered in a non-randomized strategy. In such a setting, we would need to adjust for cluster-level confounding in Stage 2 of Two-Stage TMLE. We also plan to extend this work to individually randomized trials where the health promotion strategy is delivered at the individual-level. Two-Stage TMLE, designed for hierarchical data, would not be immediately applicable in such a setting; however, we hypothesize that creating artificial clusters within arm and implementing the Two-Stage procedure (with appropriate weights) should facilitate estimation of the total effect, despite the mediation-missing data problem. Finally, in this paper as well as previous work with Counterfactual Strata Effects [9, 6, 4, 38, 21, 45], we have estimated the numerator and denominator separately prior to taking their ratio. The denominator defines being in the focus population of interest and is always bigger than the numerator. Therefore, the ratio is guaranteed to be bounded between 0 and 1. Using an efficient estimator for both the numerator and denominator implies via the Delta Method that the ratio estimator is asymptotically linear with influence function equal to the canonical gradient of the ratio parameter and, thus, achieves the semi-parametric efficiency bound [67, 59]. Nonetheless, it is of interest to develop a TMLE directly targeting the ratio. Future work is warranted.